OpenStudy (anonymous):

Last one: If 5.7 moles of zinc metal react with 8.9 moles of silver nitrate, how many moles of silver metal can be formed, and how many moles of the excess reactant will be left over when the reaction is complete? Unbalanced equation: Zn + AgNO3 → Zn(NO3)2 + Ag Be sure to show all of your work. @Nurali

OpenStudy (anonymous):

@aaronq

OpenStudy (aaronq):

what have you done so far?

OpenStudy (anonymous):

Zn + 2AgNO3 → Zn(NO3)2 + 2Ag 1 mole Zn requires 2 mole AgNO3 5.7 mole Zn requires 11.4 mole AgNO3 (5.7 x 2/1 = 11.4) AgNO3 is limiting 1 mole AgNO3 yields 1 mole Ag 8.9 moles AgNO3 yields 8.9 mole Ag 8.9 mole AgNO3 requires 4.45 mole Zn (8.9 x 1/2 = 4.45) 5.7 - 4.45 = 1.25 mole of Zn left after reaction 8.9 mole Ag can be formed 1.25 mole Zn left over after reaction is complete